determine if the ordered pair (12,4) is a solution to y=1/3x-3

Answer by jim_thompson5910(35256)

y=1/3x-3

4=(1/3)(12)-3 ... plug in (12,4)

4 = 12/3 - 3

4 = 4 - 3

4 = 1 ... which is FALSE

So (12, 4) is NOT a solution